AddressAssist

Program User Guide

First Edition, for AddressAssist Release 1.0 (Version 38.15) 26 October 1994

by Alex H. Schneider / Quarz Development

If you fail to expect the unexpected, the system will crash!

Hope, I didn’t!

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 Quarz Development / Alex H. Schneider

This is the First edition of this documentation, and it is consistent with AddressAssist Release 1.0 (Version 38.15, 26 October 1994).

Published by Quarz Development Alex H. Schneider Bachtobelstrasse 13 CH-8106 Adlikon, Switzerland, Europe

1 Introduction

AddressAssist is a brand new database program for managing addresses. But why AddressAssist??? You really might ask this question! But as always, there was one big reason for writing a new program (among many others): Many programs available in the width of the networks provided many functions, but none of them combined all my desired functions with a nice, intuitive and WB 3.1 compatible graphical user interface. This was the hour, when this program was born! It has now reached the state for its first public release with all major features included. The program is distributed as Shareware. See [Copyright and Registration], page 15, for details about your rights and the restriction of the freely distributable version.

1.1 Highlights of AddressAssist

But now, what is special with AddressAssist ? To get a short overview over the program’s possibilities, read the following list of highlight features, which are among many others included:

• 12 field address database including 4 user definable text fields.

• Very nice and intuitive user interface based on MUI (which makes it size and font sensible), 3.1 compatible

• Powerful printing interface for output address lists and labels by providing a user definable and flexible output format

• Use of address selections for printing and export

• Powerful database sort feature, possibility to use several index as well as sort in up- or downward direction

• Flexible Import/Export filters for ASCII files with user definable file format

• Import/Export filters for TransfileSF (TM) software (for transfer data with Casio pocket computer)

• Copy address data to Amiga clipboard. For use with word processors, for example.

• Easy navigation through records by a set of quick access controls (known from Superbase)

• Special function for removing identical records

• Localized under WB2.1 and above. Built-in English language, German catalog and empty catalog description file provided

• Works with WB2.04 and above

1.2 Magic User Interface (MUI)

To provide maximal comfort, the user interface of this piece of software was designed to use MUI extensively. MUI is a separate product which is copyright by Stefan Stuntz. Based on famous BOOPSI (Basic Object Oriented Programming System for Intuition), it provides a great enhancement and possibility of standardisation of the Graphical User Interface (GUI) mainly for the user. More than that, from the point of view of the developers, it makes life much easier to provide a consistent, and so called style guide conform GUI for a software project. Use MUI in your own software projects to make you programs even better!!! See ‘Readme.MUI’ for more information about MUI, its distribution and where to get it from.

2 Program overview

2.1 Program controls

The main work field of AddressAssist is the big main window, where you can recognise four different control groups:

• current record edit group with the 12 datafield string gadgets, used to edit the current address record.

• database group with the 3 buttons used for quick access to database I/O functions.

• record group with the 4 buttons used for quick access to record edit functions.

• record access control group with the 6 buttons used for quick access to a specific record in the current database.

In addition to this, there is a menu strip attached to the main window, which provides access to all available functions of the program:

• Project Menu holds project (database) related functions like file I/O and printing.

• Record Menu holds all address (record) related functions.

• Process Menu holds special functions like database import/export, sorting and the user preferences editor.

The handling of all the controls are described detailed later in this manual.

2.2 Database format

The AddressAssist’s database provides 12 record fields, which should be enough for the most address management tasks, at home and in business. In the context of this database and the manual, an ’address’ is referenced as record.

‘Salutation’

length: 20 chars; useful for printing address labels

‘Name’ length: 60 chars; first and last name or company name

‘Street1’

‘Street2’ length: 60 chars each; street and number or additional informa-

tion

‘PC/City/State’

length: 100 chars each; post code, city and state information

‘Phone1’

‘Phone2’ length: 30 chars each; useful for phone and fax numbers

‘Free1’
‘Free2’
‘Free3’

‘Free4’ length: 60 chars each; user definable fields, see below. ‘Code’ length: 10 chars; useful to distinguish between several classes of

addresses by using an appropriate character code for each of the different classes. Classes of addresses may be, for example, private or business addresses and hotels, restaurants or companies.

Using the user definable fields

If all the given fields are still not enough for you, you can define the four ‘FreeX’ fields to your own needs. Do this by giving them a name of your choice and then use them consequently in your database. The record field names can be changed in the Preferences window which is accessible through the Change preferences... item of MENU PROCESS.

2.3 Database navigation

AddressAssist provides several controls for navigating through the records of the address database. They are all accessible through the buttons of the record access control group on the bottom right of the window, but may also be reached by a keyboard shortcut: '<<' (Shortcuts: left or down arrow); moves to the previous record. '>>' (Shortcuts: right or up arrow); moves to the next record. '|<' (Shortcuts: shift left or shift down); moves to the first

record of the database.

'>|' (Shortcuts: shift right or shift up); moves to the last record

of the database.

'==' (Shortcuts: =, ramiga = or space); move to a record by entering

a key string to find it, see Section 4.1 [Find a record], page 8.

'?' (Shortcuts: ? or ramiga ?); select a record from the record table in the Select records window or create a selection of records, see Section 4.3 [Create a selection of records], page 8.

3 Database file I/O functions

The database file I/O functions are:

• Load/Save...: Load/Save in default file format (project files)

• Import/Export ASCII...: Import/Export database for various database programs.

• Import/Export Transfile...: Import/Export database for TransfileSF (TM) program.

3.1 Load/Save... project files

AddressAssist stores databases to disk by default in so called project files, which use as low disk space as possible without any compression. Project files can be read only by AddressAssist. You can access the project file I/O functions through

  • The MENU PROJECT

• Gadgets in the Database group on main window

• Shortcuts

New clears the current database in memory and you can begin to work on a new database. You will be prompted to save changes to the old database if necessary.

Load... opens the filerequester window which prompts for the filename of the address database to be loaded. Then the database will be read from disk if possible. If there is already a database in memory, it will be entirely cleared. You will be prompted to save changes to it if necessary. (Shortcuts:

o or ramiga o)

Save... writes the current database to disk with its associated filename. If there is no associated filename then you will be prompted to input a filename by the filerequester window. (Shortcut: ramiga s)

Save as... (only in Menu) opens the file requester window which prompts for the filename of the address database to be saved. (Shortcut: ramiga a)

3.2 Import/Export ASCII files

If you like to transfer the data of an AddressAssist database to or from other databases on the Amiga (f.e. Superbase, MaxonTwist) or on other platforms (f.e. MS Access, DBase, Filemaker), you may do this by using ASCII files. In AddressAssist, ASCII Files can be read/written by the Import/Export ASCII... functions in MENU PROCESS:

Import ASCII... opens the filerequester window which prompts for the filename of the ASCII data file to be imported. All records will added to the current database. Identical records can be deleted by a separate function (Remove redundant records... from MENU PROCESS) afterwards, if desirable (Section 4.4 [Remove redundant records], page 9).

Export ASCII... first opens a dialogbox where you can specify whether to export all records or only previously selected records (Section 4.3 [Create a selection of records], page 8) from the current database. Then you will be prompted to input the filename for the ASCII export data file.

The format of the ASCII data files is as follows: The end of each record is marked by a special char, the so called Record separation char and every field of the records are separated by another specific char, the Datafield separation char. The default chars are ’LF’ (ASCII 10) for record separation and ’,’ (ASCII 44) for field separation. You can change the separation chars in the Preferences window accessible through the Change preferences... item of MENU PROCESS.

3.3 Import/Export Transfile SF

If you use a Casio SF-7000, SF-7500, SF-8000, SF-9000 or SF-9500 pocket computer for your address management and also have the software package ’Transfile Amiga SF 7000’ from Yellow Computing providing the special cable for connection of the Casio to the Amiga’s serial port and the transfer program ’TFSF’, you can read or write its telephone data files (*.tele) directly with AddressAssist using the Import/Export Transfile... functions accessible from MENU PROCESS.

Import Transfile... opens the filerequester window which prompts for the filename of the Transfile telephone data file (*.tele) to be imported. All records will be added to the current database. If necessary, all identical records can be deleted by the separate function Remove redundant records... from MENU PROCESS afterwards (Section 4.4 [Remove redundant records], page 9).

Export Transfile... first opens a dialogbox where you can specify whether to export all records or only previously selected records (Section 4.3 [Create a selection of records], page 8) from the current database. Then you will be prompted to input the filename for the Transfile telephone data file.

As the Casio’s database format differs little from the one used in AddressAssist, there were some additional definitions necessary: In the Casio SF/Transfile (TFSF) database, all datafields may have more than one line entry, each line separated by a linefeed character (ASCII 10). Comparing this to the AddressAssist database, the Phone field of TFSF may have a maximum of 2 line entries (for ‘Phone1’ and ‘Phone2’ field in AddressAssist) and the Address field max. 3 line entries (the first two for ‘Street1’ and ‘Street2’ fields and the last for the ‘PC/City/State’ field). The 6 FreeX fields in TFSF are used as follows: The first four fields become the AddressAssist’s ‘Free1-4’ fields. The fifth field is the ‘Salutation’ and the sixth field becomes the ‘Code’ field in the Amiga address manager.

4 Database support functions

The database support functions are:

• Find a record: Use a key string to find a specific record.

• Sort the database: Sort database with several index and directions.

• Create a selection of records: Making selections of records for printing or exporting.

• Remove redundant records: Removing identical records from current database.

4.1 Find a record by key string

You can find a record from the current database by a key string. The program compares the ‘Name’ field of all records with the key string and returns the first record, whose ‘Name’ is matched by the key string. You can access this function from MENU RECORD or from the '==' button in the controls group on the main window.

Set key... opens a string requester prompting you to input the key string. After accepting by clicking OK, the database will examined for the given key and the first matching record is made the current one. (Shortcuts: = or

ramiga =)

4.2 Sort the database

The current database may be sorted by different index and in both ascending and descending direction. The default sort mode is ascending by index Name. You can change the sort mode through the MENU RECORD.

Sort database... opens the sort mode dialogbox, where you can choose the sort direction and select the appropriate index. By pushing the 'OK' button, the database will be resorted in the desired way. (Shortcut: ramiga

r)

The current sort mode is displayed in the title bar of the main window.

4.3 Create a selection of records

A special feature of AddressAssist is the capability of creating selections of records, which then can be printed or exported separately to all other records of the database. This function is accessible from the MENU RECORD or from the '?' button in the controls group on the main window.

Selection... opens the Select records dialogbox, which consists of a big list view gadget and three buttons for controlling the selection: All selects all records, None unselects all previously selected records and Toggle changes the selection state of the currently active record. (Shortcuts: ? or ramiga

?)

A selected record is displayed in the list view window by a white '*' in front of the name field. An unselected record has no special marking char.

4.4 Remove redundant records

A sometimes really useful feature is the possibility to remove identical records from the current database. For example, if you use several database programs on different platforms (at home and at work) or the Casio SF pocket computer for address management: if you then like to put together all records from two ore more databases, you would have to remove all the identical records manually from the joined database. In AddressAssist this can be done with the Remove redundant records... function, which is accessible

from MENU RECORD.

Remove redundant records... examines the database for identical records and prompts whether or not to delete found identical records. When you specify to delete All, you will not be prompted anymore. There will be a message at the end of the operation confirming, that there are no more redundant records in the database memory.

5 Database print function

AddressAssist provides a really powerful print interface which enables you to print your addresses with many (user definable) print formats for address lists or labels. You may specify all relevant parameters for text printing including, for example, how many records should be printed on one sheet of paper (very useful for label printing) or whether or not to print the record field names. For each datafield, you furthermore may specify separately whether or not it should be printed at all, which style should be used for printing and how many linefeeds should be printed after each field. All these parameters can be changed in the Print Setup dialogbox, which is accessible through the Print... item of MENU PROJECT. The dialogbox consists of three pages, which are described separately:

• Page setup: # linefeed before/after record. # records per page.

• Record field options: Select datafield, style and format.

• Printer setup: Output device: printer or file; paper length and margins.

If you choose 'OK', the program will print with the specified parameters. By selecting 'Save', you are able to store recently changes to the print parameters without printing. Leave the dialogbox with 'Cancel'.

5.1 Page setup

In the Page setup page, you are able to change the page format parameters to your desire and select, whether to print all records or only previously selected records.

Print Mode toggles between printing or not the datafield names. This is useful while printing labels, where you don’t like field names as against of printing an address list.

Selection toggles between printing all or only previously selected records. Section 4.3 [Create a selection of records], page 8.

LF before address and LF after address indicate the number of linefeeds to print in front of / after the records.

With Records per page it is very easy to take care of not to split addresses on two pages or to print the desired number of addresses on a predefined label sheet. Use this together with Page Length from the Print setup page. Set Records per page to a zero or a value smaller than zero to prevent AddressAssist from checking the number of addresses printed on one sheet of paper.

5.2 Record field options

For each record field, there are 3 separate controls available, which make even the craziest formats for printing a record possible:

Use the Checkbox to indicate whether or not to print this field at all.

The Linefeed Cycle changes the number of linefeed to be printed after each record field. You can select from no LF up to 3 LF.

The Style Cycle changes the print style for the record field. You can choose

between Normal, Bold, Italic or Bold-Italic.

5.3 Printer setup

In the Printer setup page, you may select the output device: output to printer or output into an ASCII text file. When using the printer for output, you furthermore may change current printer page length and paper margins.

The Printer text field shows the currently selected preferences printer driver. Change the printer driver by using the Printer preferences program from your workbench distribution.

Page Length indicates the number of lines, which can be printed on one sheet of paper.

Top Margin is the number of empty lines on the top of the paper.

The cycle gadget Line Spacing toggles between six or eight lines per

inch.

Left Margin indicates the left margin on the paper in number of characters.

Right Margin indicates the right margin on the paper in characters, counted from the left side.

Output to toggles current print output device: Printer (PRT:) or text File.

6 Record support functions

The record support functions include all available functions for working on and with individual records:

• New record: Create a new record.

• Save record: Save current record.

• Edit record: Start to edit the current record.

• Duplicate record: Duplicate the current record.

• Delete record: Remove the current record from database.

• Copy record: Copy current record to clipboard.

6.1 New record function

The new record function is accessible through the MENU RECORD or through the 'New' button in the record group of the main window.

New cleans the display of the current record, after prompting you to save changes to the cleaned record if appropriate, and then activates the first string gadget for input of a new record. (Shortcuts: n or ramiga n)

6.2 Save record function

The save record function is accessible through the MENU RECORD or through the 'Save' button in the record group of the main window (Shortcut s). Additionally you will be prompted automatically for the save function, either when leaving the last datafield string gadget (‘Code’) by a return or by selecting a function which would cause the loss of recently changed data.

Save saves recent changes to the current record or adds a the current as a new record to the database. (Shortcut: s)

6.3 Edit record function

The edit record function is accessible through the MENU RECORD.

Edit activates the first datafield of the current record (‘Salutation’), which is the same action as when clicking with the mouse on this gadget or another of current record edit group’s string gadgets. This is useful, if the program is used only by keyboard. (Shortcut: e or ramiga e)

6.4 Duplicate record function

The duplicate record function is accessible through the MENU RECORD or through the 'Duplicate' button in the record group of the main window.

Duplicate adds a new record to the database using the data of the current record. You will be prompted by a requester to confirm this action. (Short-

cut: d or ramiga d)

6.5 Delete record function

The delete record function is accessible through the MENU RECORD or through the 'Delete' button in the record group of the main window.

Delete removes the current record from memory. You will be prompted by a requester to confirm this unrecoverable action. (Shortcut: l or ramiga l)

6.6 Copy record function

The copy record function is accessible only through the MENU RECORD.

Copy copies the text data of the current record to the Amiga’s clipboard, where it can be read and used from other application. Use this to copy the current address of AddressAssist into to a word processor or a text editor. (Shortcut: ramiga c)

7 Program history & future developments

7.1 Program history

• Version 1.0 (26.10.94), version 38.15:

- First public release.

7.2 Future developments

This is a list of possible future enhancements for AddressAssist. Remember, the future of this program also depends on you, the user: register immediately and make suggestions, bug reports.

• Enhance Selection interface:

- Implementing Filter function

- Load/Save selections

• Enhance printer interface:

- Adding database field length for line list printing

- Adding header and footer line feature

- Load/Save individual print setup

• Enhance Import/Export of Transfile/Casio SF data: - Add direct read/write filters for Casio SF from serial device. (Does anybody know how to write data to Casio SF pocket computers, reading is no problem...)

• Add application specific functions to the AREXX Port

• Add address grouping feature

• Add german documentation (sorry for now!!!)

8 Copyright and Registration

This program, AddressAssist, is distributed as Shareware. The freely distributable archive contains the unregistered version of the program. It is limited to work with only 15 records. The freely distributable archive may be freely redistributed (as its name says!), but no charges other than reasonable copying and handling fees may be collected. The program may be distributed only as originally released, in the complete archived form. No changes of any kind may be made to the AddressAssist binary, documents or archive, and I will not answer for any problems that may result from a modified binary. The suggested basic registration fee is SFr 30 (Swiss Francs), DM 30 or US$ 25 (no other currencies accepted), which gets you an entry in my database of users and a certificate of registration by return mail. You will get the latest release of the program together with your personal key file. This key file disables the restrictions of 15 records and will work with all updates of the program that will be available on the networks in the future. This key file is personal and definitely not freely distributable. It may be used only for one application of AddressAssist at the time. If you like to run many copies of AddressAssist in the same time, please ask for detailed information about licensing. If you like to register now, see also file ‘Orderform.txt’.

Trademarks

Transfile Amiga SF 7000 is a registered trademark of Yellow Computing. Amiga, Amiga Workbench and Amiga Kickstart are registered trademarks of Commodore-Amiga Inc. Installer and Installer project icon are \begin{matrix}c \\ ◯ \end{matrix} Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved. MUI is \begin{matrix}c \\ ◯ \end{matrix} Copyright 1993-94 Stefan Stuntz (see ‘Readme.MUI’) for details. All other named trademarks reference to their appropriate manufactor or seller.

9 Disclaimer

THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDER AND/OR OTHER PARTIES PROVIDE THE PROGRAM "AS IS" WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MER-CHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFEC-TIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSE-QUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

Chapter 10: Contacting the author 17

10 Contacting the author

For questions, registration, suggestions or bug reports please contact me by mail: Alex H. Schneider Bachtobelstrasse 13 CH-8106 Adlikon Switzerland / Europe or by email, internet (no registration): schneidr@igt.baum.ethz.ch schneida@hoggar-ether.ethz.ch or by phone (no registration): ++41 1 841 10 17 (at home) ++41 1 633 30 40 (at work)

Index 18

Index

C

Changing of the FreeX fields. . . . . . . . . . . . 3 Contacting the author. . . . . . . . . . . . . . . . . 17 Copy record function . . . . . . . . . . . . . . . . . . 13 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Create a selection of records . . . . . . . . . . . . 8 Current record edit group . . . . . . . . . . . . . . 3

D

Database file I/O functions . . . . . . . . . . . . . 5 Database format . . . . . . . . . . . . . . . . . . . . . . . 3 Database group . . . . . . . . . . . . . . . . . . . . . . . . 3 Database print function . . . . . . . . . . . . . . . 10 Database support functions. . . . . . . . . . . . . 8 Delete record function . . . . . . . . . . . . . . . . . 13 Duplicate record function . . . . . . . . . . . . . 13

E

Edit record function. . . . . . . . . . . . . . . . . . . 12

F

Find a record by key string . . . . . . . . . . . . . 8 Future . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

H

History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

I

Import/Export ASCII files . . . . . . . . . . . . . 5 Import/Export Transfile SF . . . . . . . . . . . . 6 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

L

Load/Save... project files . . . . . . . . . . . . . . . 5

N

New record function. . . . . . . . . . . . . . . . . . . 12

P

Page setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Printer setup. . . . . . . . . . . . . . . . . . . . . . . . . . 10 Program controls . . . . . . . . . . . . . . . . . . . . . . . 3

R

Record access control group . . . . . . . . . . . . 3 Record field options (printing). . . . . . . . . 10 Record group . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Record support functions. . . . . . . . . . . . . . 12 Registration. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Remove redundant records . . . . . . . . . . . . . 9

S

Save record function . . . . . . . . . . . . . . . . . . 12 Sort the database . . . . . . . . . . . . . . . . . . . . . . 8

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Highlights of AddressAssist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Magic User Interface (MUI) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Program overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Program controls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Database format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Using the user definable fields . . . . . . . . . . . . . . . . . . . . 4

2.3 Database navigation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Database file I/O functions . . . . . . . . . . . . . . . . . . . . 5

3.1 Load/Save... project files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Import/Export ASCII files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Import/Export Transfile SF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4 Database support functions . . . . . . . . . . . . . . . . . . . . 8

4.1 Find a record by key string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 Sort the database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.3 Create a selection of records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.4 Remove redundant records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

5 Database print function. . . . . . . . . . . . . . . . . . . . . . . 10

5.1 Page setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.2 Record field options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.3 Printer setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6 Record support functions . . . . . . . . . . . . . . . . . . . . . 12

6.1 New record function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.2 Save record function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.3 Edit record function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.4 Duplicate record function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.5 Delete record function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.6 Copy record function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

7 Program history & future developments …. 14

7.1 Program history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 7.2 Future developments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

8 Copyright and Registration. . . . . . . . . . . . . . . . . . . 15

Trademarks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

9 Disclaimer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

10 Contacting the author . . . . . . . . . . . . . . . . . . . . . . . 17

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18